# Conformal Prediction Algorithms for Time Series Forecasting: Methods and Benchmark

### v1

@@ Abstract
Reliable uncertainty quantification is of critical importance in time series forecasting, yet traditional methods often rely on restrictive distributional assumptions. Conformal prediction (CP) has emerged as a promising distribution-free framework for generating prediction intervals with rigorous theoretical guarantees. However, applying CP to sequential data presents a primary challenge: the temporal dependencies inherent in time series fundamentally violate the core assumption of data exchangeability, upon which standard CP guarantees are built. This review critically examines the main categories of algorithmic solutions designed to address this conflict. We survey and benchmark methods that relax the exchangeability assumption, those that redefine the data unit to be a collection of independent time series, approaches that explicitly model the dynamics of the prediction residuals, and online learning algorithms that adapt to distribution shifts to maintain long-run coverage. By synthesizing these approaches, we highlight computational efficiency and practical performance on real-world data.
@@ 1 Introduction
2 Foundational Concepts: Split Conformal Prediction
@@ 3 Conformal Prediction Algorithms for Time Series
3.1 Split Conformal Prediction Adapted to Multi-Horizon Forecasting (MSCP)
Core principle. Standard SCP can be adapted to multi-step split conformal prediction (MSCP) by constructing horizon-specific calibration scores and quantiles (Wang and Hyndman, 2024, 2026). Unlike standard CP, which often relies on random splitting, time-series data requires a sequential split to preserve the underlying temporal dependencies and prevent data leakage.
@@ 3.2 Ensemble Batch Prediction Intervals (EnbPI)
Methodology. Train an ensemble of $B$ base models on bootstrap resamples of the full training set. For each training point $i$, form a leave-one-out (LOO) aggregate predictor $\tilde{y}_{i}$ by averaging only those ensemble members not trained on point $i$. The LOO residuals $\tilde{r}_{i}=|y_{i}-\tilde{y}_{i}|$ serve as calibration scores, leveraging all $T$ points without a strict train/calibration split. As data arrives, maintain a sliding window of residuals to adapt to recent dynamics.
@@ 3.4 Conformal Prediction for Time Series Dataset (Global-CP)
3.5 Online Learning and Control (ACI, AcMCP)
@@ 4.2 Evaluation Metrics
Winkler interval score: Composite metric penalizing both miscoverage and unnecessary width. For observation $i$ with bounds $(\ell_{i},u_{i})$:
@@ 4.3 Benchmark Results
Figure 1: Empirical coverage rate (dashed orange line represents target coverage).
@@
Figure 3: Critical difference diagram based on the Winkler interval score for top 3 algorithms
@@ 5 Conclusion
In conclusion, this study highlights that for practitioners utilizing modular forecasting pipelines, the MSCP approach offers the most robust and efficient solution for distribution-free uncertainty quantification. The success of this method emphasizes the importance of horizon-specific calibration in accounting for the temporal dependencies that typically violate the exchangeability assumption. While online controllers like ACI and AcMCP remain highly competitive and adaptable to non-stationary environments, the simplicity and superior efficiency of horizon-specific split conformal prediction make it the preferred choice for the high-stakes sequential tasks evaluated here.

### v2

@@ Abstract
Reliable uncertainty quantification is of critical importance in time series forecasting, yet traditional methods often rely on restrictive distributional assumptions. Conformal prediction (CP) has emerged as a promising distribution-free framework for generating prediction intervals with rigorous theoretical guarantees. However, applying CP to sequential data presents a primary challenge: the temporal dependencies inherent in time series fundamentally violate the core assumption of data exchangeability, upon which standard CP guarantees are built. This paper critically examines the main categories of algorithmic solutions designed to address this conflict. We survey and benchmark methods that relax the exchangeability assumption, those that redefine the data unit to be a collection of independent time series, approaches that explicitly model the dynamics of the prediction residuals, and online learning algorithms that adapt to distribution shifts to maintain long-run coverage. We use AutoARIMA as the base forecaster on a large-scale monthly sales dataset, evaluating marginal coverage, interval width, and the Winkler score. Our benchmark results show that multi-step split conformal prediction method meets the 90% coverage threshold and demonstrates the best efficiency.
@@ 1 Introduction
2 Foundational Concepts
@@ 3 Conformal Prediction Algorithms for Time Series
3.1 Multi-step Split Conformal Prediction (MSCP)
Core principle. MSCP adapts standard SCP to the multi-horizon forecasting setting by constructing horizon-specific calibration scores and quantiles (Wang and Hyndman, 2024, 2026). Unlike standard CP, which often relies on random splitting, time-series data requires a sequential split to preserve the underlying temporal dependencies and prevent data leakage.
@@ 3.2 Ensemble Batch Prediction Intervals (EnbPI)
Methodology. Train an ensemble of $B$ base models on bootstrap resamples of the full training set. For each training point $i$, form a leave-one-out (LOO) aggregate predictor $\tilde{y}_{i}$ by averaging only those ensemble members not trained on point $i$. The LOO residuals $\tilde{r}_{i}=|y_{i}-\tilde{y}_{i}|$ serve as calibration scores, leveraging all $T$ points without a strict train/calibration split. As data arrive, maintain a sliding window of residuals to adapt to recent dynamics.
@@ 3.4 Conformal Prediction for Time Series Dataset (Global-CP)
3.5 Online Learning and Control
@@ 4.2 Evaluation Metrics
Winkler interval score (WIS): Composite metric penalizing both miscoverage and unnecessary width. For observation $i$ with bounds $(\ell_{i},u_{i})$:
@@ 4.3 Benchmark Results
Figure 1: Empirical coverage rate (dashed orange line represents target coverage of 90%).
@@
Figure 3: Critical difference diagram based on the Winkler interval score for the top 3 algorithms: MSCP, Parametric-PI, and ACI
@@ 5 Conclusion
In conclusion, this study highlights that for practitioners utilizing modular forecasting pipelines, the MSCP approach offers the most robust and efficient solution for distribution-free uncertainty quantification in multi-horizon forecasting. The success of this method emphasizes the importance of horizon-specific calibration in accounting for the temporal dependencies that typically violate the exchangeability assumption. While online controllers like ACI and AcMCP remain highly competitive and adaptable to non-stationary environments, the simplicity and superior efficiency of horizon-specific split conformal prediction make it the preferred choice for the high-stakes sequential tasks evaluated here.